OpenStudy (anonymous):

i am having a problem with the square root of 5x^2-3x=2x

OpenStudy (anonymous):

square root inside the quadratic formula?

OpenStudy (anonymous):

The square root of...? The roots are 0 and 1.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Can you rephrase your question? I'm not quite sure what's being asked.

OpenStudy (anonymous):

I do not think we want to take the square root of 5x^2-3x=2x??

OpenStudy (anonymous):

ii have to solve the rational expone t

OpenStudy (anonymous):

I am still unclear what is being asked, but if I was given 5x^2-3x=2x with no directions, I would subtract 2x from both sides 5x^2 - 5x = 0 then factor out the GCF 5x(x - 1) = 0 and solve, with the solutions being x = 0 and x = 1.

OpenStudy (anonymous):

Yeah I'm not quite sure either. A square root is a rational exponent: \[\Large \sqrt x = {x^{1/2}}\] But I still don't really know what you're asking zonk...?

OpenStudy (anonymous):

thats what i got to but which is the answer or are both the answer

OpenStudy (anonymous):

Can you explain what you're actually doing zonk, and show the steps you've taken?